Some say Hillary Clinton is unstoppable for 2016. A recent poll indicates that 586/1000 voters will favor her in the next election.

A) Create a 90% confidence interval for P ( the population proportion) and interpret.

B) Based off of that interval is there strong evidence to say she will win?

C) How large of a sample is needed for a desired margin of error of 2% using 90% confidence and 586/1000 as a point estimate for P.
1
Expert's answer
2020-04-17T13:27:07-0400

A) We need to construct the 90% confidence interval for the population proportion. We have been provided with the following information about the number of favorable cases:

"Favorable\ Cases \ X=586"

"Sample\ Size \ N=1000"

The sample proportion is computed as follows, based on the sample size "N=1000" and the number of favorable cases"X=586."


"\hat{p}={X \over N}={586 \over 1000}=0.586"

The critical value for "\alpha=0.1"  is "z_c=z_{1-\alpha/2}=1.645." The corresponding confidence interval is computed as shown below:


"CI(Proportion)=\big(\hat{p}-z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}},\hat{p}+z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}}\big)"

"=\big(0.586-1.645\sqrt{{0.586(1-0.586) \over 1000}},0.586+1.645\sqrt{{0.586(1-0.586) \over 1000}}\big)"

"=(0.560378,0.611622)"

Therefore, based on the data provided, the 90% confidence interval for the population proportion is "0.560378<p<0.611622," which indicates that we are 90% confident that the true population proportion "p" is contained by the interval "(0.560378,0.611622)."


B) Based off of that interval there is strong evidence to say she will win "(p>0.5)."


C) Margin of error "E"


"E=z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}}\leq0.02"

"N\geq{(1.645)^2 (0.586)(1-0.586) \over (0.02)^2}"

"N\geq1642"

"1642"